OpenStudy (anonymous):

what is the length of DK.....the answer choices are A. 8.5 B. 7.0 C. 12.7 D. 3.5

OpenStudy (igreen):

Picture?

OpenStudy (anonymous):

idk how to send you the pic..but its a circle and at the side of the circle is AK= 14 EK=17 B=7

OpenStudy (igreen):

Save it to your computer and use the Attach File button, or right click and copy image url then paste it here.

OpenStudy (phi):

the products of the segments are equal

OpenStudy (anonymous):

im horrible at this.. I need you to explain it to me please

OpenStudy (phi):

Do you see the "chord" BE ? It is divided into two parts (by point K) what are the lengths of each part?

OpenStudy (anonymous):

each part of what chords?

OpenStudy (phi):

do you see BE ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

do you see point K on BE ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

so BE is divided into two segments by point K what are the lengths of each of those segments?

OpenStudy (anonymous):

BK is 7

OpenStudy (phi):

and the other one?

OpenStudy (anonymous):

idk be

OpenStudy (phi):

you want the length fro K to E

OpenStudy (phi):

*from

OpenStudy (phi):

it's the same as the length from E to K

OpenStudy (anonymous):

EK is 17

OpenStudy (phi):

yes (and hopefully it's clear that KE is the same length) now you have the two parts of chord BE you need the product of the two lengths (that means multiply the two lengths)

OpenStudy (anonymous):

so multiply ke and be together ?

OpenStudy (phi):

ke and kb (not be)

OpenStudy (anonymous):

so 17 and 17?

OpenStudy (phi):

ke is 17 kb is 7

OpenStudy (anonymous):

its 119

OpenStudy (phi):

ok, now we do the other chord , chord AD what is the length of each segment of chord AD ?

OpenStudy (anonymous):

AK is 14

OpenStudy (phi):

and what is the other segment?

OpenStudy (phi):

it is what we have to find: what is the length of DK.. I would call it length "x" (x for unknown) now find the product of 14 and x do you know how to write 14 times x?

OpenStudy (anonymous):

yes

OpenStudy (phi):

how do you write it ?

OpenStudy (anonymous):

14x

OpenStudy (phi):

and the idea is those two products are equal in other words 14x = 119

OpenStudy (phi):

to find x, divide both sides by 14 \[ \frac{14}{14}x = \frac{119}{14} \]

OpenStudy (anonymous):

8.5

OpenStudy (phi):

on the left side 14/14 is 1, so the left side is just x yes, you get x= 8.5

OpenStudy (anonymous):

i got it now! thank you!!

OpenStudy (phi):

If I were doing it I would write \[ 14x= 7\cdot 17\] then divide both sides by 14 to get \[ x = \frac{7\cdot 17}{14} \] I know 7/14 is 1/2 so this is \[ x= \frac{17}{2} \] and finally x= 8.5